Section 06  Topology

Section 06: Topology

Invited Lectures

Wednesday, 19 August  MA 004

14:00 – 14:45 Robert A. Oliver, Université Paris-Nord, Villetaneuse, France Vector bundles over classifying spaces (I: p.84) Chair: Charles B. Thomas

15:00 – 15:45 Ronald A. Fintushel, Michigan State University, East Lansing, USA Ronald J. Stern, University of California, Irvine, USA Knot Surgery on 4-Manifolds (I: p.86) Chair: Ulrich Koschorke

Thursday, 20 August  MA 004

16:15 – 17:00 Tomotada Ohtsuki, Tokyo Institute of Technology, Tokyo, Japan A filtration of the set of integral homology 3-spheres (I: p.82) Chair: W.B. Raymond Lickorish

17:15 – 18:00 Clifford Henry Taubes, Harvard University, Cambridge, USA The Geometry of the Seiberg-Witten Invariants (I: p.86) Chair: Dieter Kotschick

Friday, 21 August  MA 004

14:00 – 14:45 Alexander N. Dranishnikov, University of Florida, Gainesville, USA and Steklov Math. Institute, Moscow, Russia Dimension theory and large Riemannian manifolds (I: p.79) Chair: Alexander S. Mishchenko

15:00 – 15:45 Michael Hartley Freedman, Microsoft Corporation, Redmond, USA Topological views of computational complexity (I: p.80) Chair: Robert D. Edwards

Saturday, 22 August  MA 004

16:15 – 17:00 William Gerard Dwyer, University of Notre Dame, Notre Dame, USA Lie groups and p-compact groups (I: p.79) Chair: Albrecht Dold

17:15 – 18:00 Mark Mahowald, Northwestern University, Evanston, USA Toward a global understanding of π_{∗}(S^{n}) (I: p.81) Chair: Albrecht Dold

Topology  Section 06

Short Communications and Poster Sessions

Short Communications  Wednesday, 19 August
Group Actions  H 112

Chair: Reinhard E. Schultz

16:15 – 16:30 Charles B. Thomas, University of Cambridge, Cambridge, Great Britain Contact structure and spherical space forms (p.98)

16:30 – 16:45 Harald Biller, Universität Stuttgart, Stuttgart, Germany Actions of large compact Lie groups on spheres (p.85)

16:45 – 17:00 Takashi Tsuboi, University of Tokyo, Tokyo, Japan On the group of area preserving homeomorphisms (p.99)

Knots  H 112

Chair: Hitoshi Murakami

17:15 – 17:30 Alan H. Durfee, Mount Holyoke College, South Hadley, MA, USA Polynomial representations of knots (p.87)

17:30 – 17:45 Hugh R. Morton, University of Liverpool, Liverpool, Great Britain The Burau matrix and Fiedler’s invariant for closed braids (p.94)

17:45 – 18:00 Vladimir V. Vershinin, Sobolev Institute of Mathematics, Novosibirsk, Russia Generalizations of Braids: Homology and Vassiliev Invariants (p.100)

Short Communications  Thursday, 20 August
Gauge Theory and Low Dimensional Topology  H 112

Chair: Ronald J. Stern

14:00 – 14:15 Yong S. Cho, Ewha Women’s University, Seoul, Korea, Republic of Finite Group Actions and Seiberg-Witten Invariants (p.87)

14:15 – 14:30 Nikolai Saveliev, University of Michigan, Ann Arbor, MI, USA Casson invariant and surface singularities (p.97)

14:30 – 14:45 Louis H. Kauffmann, University of Illinois, Chicago, USA Functional Integrals and the Kontsevich Integral (p.90)

Knots and 3-Manifolds  H 112

Chair: Jonathan A. Hillman

15:00 – 15:15 Joel Hass, University of California, Davis, CA, USA The number of Reidemeister Moves Needed for Unknotting (p.89)

15:15 – 15:30 Cherry Kearton, University of Durham, Durham, Great Britain Alexander ideals of classical knots (p.90)

15:30 – 15:45 Michelangelo Vaccaro, Roma, Italy Unicity of the topological type of the simply connected 3-manifolds (p.99)

Short Communications  Friday, 21 August
Geometric Topology  H 112

Chair: Alexander N. Dranishnikov

16:15 – 16:30 Dusan Repovs, University of Ljubljana, Ljubljana, Slovenia On Banach-Mazur compacta and their geometric properties (p.96)

16:30 – 16:45 Washington Mio, Florida State University, Tallahassee, FL, USA On the geometric topology of homology manifolds (p.94)

16:45 – 17:00 Arkady Skopenkov, Dept. of Mathematics, Kolmogorov College, Moscow, Russia A Generalization of The Haefliger-Weber Theorem on Embeddings of Manifolds in ℝ^{m} (p.97)

Hyperbolic 3-Manifolds  H 112

Chair: Nikolai Gusevskii

17:15 – 17:30 Vitalii S. Makarov, Univ. Drujb. Narod., Moscow, Russia Regular maps and hyperbolic 3-manifolds (p.92)

17:30 – 17:45 Alexander D. Mednykh, Sobolev Institute of Mathematics, Novosibirsk, Russia The volumes and isometries of Heegaard genus two hyperbolic three manifolds (p.93)

17:45 – 18:00 Andrei Yu. Vesnin, Sobolev Institute of Mathematics, Novosibirsk, Russia On 3-manifolds with cyclic symmetry (p.100)

Short Communications  Saturday, 22 August
Algebraic Topology I  H 112

Chair: Jean-Michel Lemaire

14:00 – 14:15 Igor Kriz, University of Michigan, Ann Arbor, Michigan, U.S.A. Real-oriented homotopy theory (p.91)

14:15 – 14:30 Carlos Prieto, Instituto de Matematicas, Mexico, D.F., Mexico Quasifibrations and Bott periodicity (p.95)

14:30 – 14:45 Alexander S. Mishchenko, Moscow State University, Moscow, Russia Local combinatorial Hirzebruch formula (p.94)

Algebraic Topology II  H 112

Chair: Daniel H. Gottlieb

15:00 – 15:15 Carles Casacuberta, Universitat Autònoma de Barcelona, Bellaterra, Spain A universal space for plus-constructions (p.86)

15:15 – 15:30 Samson Saneblidze, A. Razmadze Mathematical Institute, Tbilisi, Georgia Adams Cobar Construction as a Cubical Complex (p.97)

15:30 – 15:45 Hirotaka Tamanoi, University of California Santa Cruz, Santa Cruz, USA BP cohomology theory and v_{n}-series (p.98)

Short Communications  Monday, 24 August
Manifolds and Dynamics  H 112

Chair: Tadayoshi Mizutani

16:15 – 16:30 Levan A. Alania, Moscow St. Univ., Dept. of Mech. & Math, Moscow, Russia De Rham cohomology with local coefficients (p.85)

16:30 – 16:45 Alexander Fel’shtyn, E.-M.-Arndt-Universität Greifswald, Greifswald, Germany Reidemeister torsion and dynamical zeta functions in Nielsen theory (p.88)

16:45 – 17:00 Efthimios Kappos, University of Sheffield, Sheffield, England Manifolds: Differential Topology: Controllability of Vector Fields (p.90)

4-Manifolds  H 112

Chair: Alexander D. Mednykh

17:15 – 17:30 J. Scott Carter, University South Alabama, Mobile, AL, United States of America The Diagrammatics of Hopf Categories and Invariants of 4-Dimensional Manifolds (p.86)

17:30 – 17:45 Nikolai Gusevskii, Universidade Federal de Minas Gerais, Belo Horizonte, Brazil Complex hyperbolic surfaces and discrete groups in complex hyperbolic geometry (p.88)

17:45 – 18:00 Alexander Prishlyak, Taras Shevchenko Kiev University, Kiev, Ukraine Morse function conjugation on 4-manifolds (p.95)

Topology  Section 06

Short Communications  Tuesday, 25 August
2-Manifolds and Related Questions  H 112

Chair: Carl-Friedrich Bödigheimer

14:00 – 14:15 José-Luis Estévez, Universidad Nacional de Educación a Distancia, Madrid, Spain A presentation for the mapping class group of a non-rientable surface from Hatcher and Thurston’s

(p.87)

14:15 – 14:30 Dmitri V. Millionshikov, Moscow State University, Moscow, Russia Krichever-Novikov algebras (p.93)

14:30 – 14:45 Andreas Zastrow, Ruhr-Universität Bochum, Bochum, Germany Planar Sets are Aspherical (p.100)

Algebraic Topology III  H 112

Chair: Leonard Mdzinarishvili

15:00 – 15:15 Ivan Ivansic, Faculty of Electrical Engeneering, University of Zagreb, Zagreb, Croatia Borsuk’s Index and Pointed Movability for Projective Movable Continua (p.89)

15:15 – 15:30 Sibe Mardesic, Department of Mathematics, University of Zagreb, Zagreb, Croatia Strong shape for topological spaces (p.92)

15:30 – 15:45 Mihkail M. Postnikov, Steklov Math. Inst., Moscow, Russian Federation Totality versus naturality (p.95)

Short Communications  Wednesday, 26 August
Algebraic Topology of Manifolds  H 112

Chair: Lawrence W. Conlon

14:00 – 14:15 Ulrich Koschorke, Universität-Gesamthochschule Siegen, Siegen, Germany μ-invariants of link maps in arbitrary manifolds (p.90)

14:15 – 14:30 Yuri V. Muranov, Vladimir State University, Vladimir, Russia Surgery and splitting obstruction groups for one-sided submanifolds (p.94)

14:30 – 14:45 Gerard A. Venema, Calvin College, Grand Rapide, Michigan, USA Representing homology classes of 4-manifolds (p.99)

Lusternik-Shnirelmann Category  H 112

Chair: Hans Scheerer

15:00 – 15:15 Jean-Michel Lemaire, CNRS/SPM, Paris, France Rational Lusternik-Schnirelmann category of Poincaré complexes (p.91)

15:15 – 15:30 Yuli B. Rudyak, Universität-GH Siegen, Siegen, Germany Lusternik–Schnirelmann category of symplectic manifolds and the Arnold conjecture (p.96)

Poster Session  Monday, 24 August 14:00 – 15:45  MA 144

Laura M. Anderson, Texas A&M University, College Station, TX, USA

Matroid bundles and characteristic classes (p.85) Lorena Armas-Sanabria, Instituto de Matematicas, UNAM, Mexico D.F., Mexico

3-Pure Braids and Surgery (p.85) Vitali Chatyrko, Linköping Univertitet, Linköping, Sweden

On infinite-dimensional Cantor manifolds (p.86) Mario Eudave-Munoz, Instituto de Matematicas, UNAM, Mexico, D.F., Mexico

Incompressible surfaces in the exterior of tunnel number one knots (p.88) Norikazu Hashiguchi, Nihon University, Tokyo, Japan

The surgery formula for the discrete Godbillon-Vey invariant (p.89) Thang T.Q. Le, SUNY Buffalo, Buffalo, NY, USA

On Quantum and Finite Type Invariants of Homology 3-Spheres (p.91) Jouni Luukkainen, University of Helsinki, Helsinki, Finland

The fundamentals of Lipschitz and quasiconformal manifolds (p.92) Leonard Mdzinarishvili, Georgian Technical University, Tbilisi, Georgia

Partially continuous Alexander-Spanier cohomology theory (p.93) Alexey Rasskazov, Novosibirsk State University, Novosibirsk, Russia

The structure of the canonical fundamental set for the 2-bridge link (p.95) Sascha A. Rogmann, Heinrich-Heine-Universität Düsseldorf, Uedemerbruch, Germany

Growth functions of tessellations (p.96) Richard J. Steiner, University of Glasgow, Glasgow, Great Britain

Pasting in multiple categories (p.98) Paul R. Turner, University of Aberdeen, Aberdeen, Great Britain

The homology of spaces representing exact cohomology theories (p.99)